In Problem 64, you’re essentially solving a problem like Example 9 (p 556). Use one ordered pair to find k, and then use that value along with the other given value to determine what its corresponding ordered pair will be.
64) Business and finance. The revenue for a sandwich shop is directly proportional to its advertising budget. When the owner spent $2000 a month on advertising, the revenue was $120,000. If the revenue is now $180,000, how much is the owner spending on advertising?






Answer by tutorcecilia(2144)

Use the formula for direct variation:
y=xk [plug-in the values]
$120,000 = $2000k [solve for the k-term]
120,000/2000=k
60=k
So, the formula becomes:
y=60x
.
If the revenue is now $180,000, how much is the owner spending on advertising?
Using y=60x [plug-in the values and solve for advertisment (x)]
($180,000) = 60x
180,000/60=60/x
$3,000 = x (amont spent on advertisment)